OpenStudy (anonymous):

What is the partial sum for the geometric series in which a3=4, a4=-16, and n=6

OpenStudy (campbell_st):

well do you know the common ratio..? if not its \[\frac{a_{4}}{a_{3}} = \]

OpenStudy (anonymous):

-16/4=-4

OpenStudy (campbell_st):

ok thats great now you need to find the 1st term so the general term is \[a_{n} = a \times r^{n -1}\] use the 3rd term... or the 4th term so for the 3rd term = 4 and common ratio r = -4 and n = 3 you have \[4 = a \times (-4)^{3 -1} \] can you find the value of a from the equation above?

OpenStudy (anonymous):

a=1/4 ?

OpenStudy (campbell_st):

correct... all you need now is the partial sum, use the formula the formula is \[s_{n} = \frac{a(1 - r^n)}{1 - r}\] you have a = 1/4, r = -4 and n = 6 substituting you get \[s_{6} = \frac{\frac{1}{4}(1 - (-4)^6)}{1 - (-4)}\] just evaluate

OpenStudy (campbell_st):

hope this helps...

OpenStudy (anonymous):

i got 204.85

OpenStudy (campbell_st):

well I got - 204.75 so me the sum should be negative since the 6th term is - 256

OpenStudy (campbell_st):

so you are adding 1/4 + -1 + 4 + -16 + 64 + -256 = hope this helps.

OpenStudy (anonymous):

ok thank you!